OpenStudy (anonymous):

How would you solve log base 4 of 12 with a calculator?

OpenStudy (anonymous):

unfortunately both answers above are incorrect. should be the log of the input divided by the log of the base

OpenStudy (anonymous):

\[\log_b(x)=\frac{\log(x)}{\log(b)}\] so in your case it should be \[\log_4(12)=\frac{\log(12)}{\log(4)}\]